OpenStudy (anonymous):

multiple to find the exact answer your answer should be a whole number, fraction or improper fraction only 6 1/3 x 1 5/8

OpenStudy (anonymous):

38?

OpenStudy (mathmale):

Multiply. I'd suggest you change the mixed number 6 1/3 into an improper fraction first: Think of 6 1/3 as \[6\frac{ 1 }{ 3 }=6+\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

but how did you get the answer explain please

OpenStudy (mathmale):

Combine these two terms to obtain a fraction. Then do the same thing to the other number (1 5/8). Then multiply the two fractions together.

OpenStudy (anonymous):

im getting 19/3 and 13/3

OpenStudy (anonymous):

is this multiple answers like a b c d ?

OpenStudy (anonymous):

nooooo

OpenStudy (mathmale):

Note well, @gentleman313, that 6 1/3 is approx. 6, and that 1 5/8 is approx. 2, so there's no way that the answer could be 38. Please refrain from passing out answers. Instead, guide the student towards finding his or her own answer.

OpenStudy (mathmale):

im getting 19/3 and 13/3 ... right. Now multiply these fractions together.

OpenStudy (anonymous):

ok then i get 247/9 which i feel is wrong

OpenStudy (anonymous):

well is it 10?? since wat ur saying?

OpenStudy (mathmale):

What is 13*19?

OpenStudy (anonymous):

247

OpenStudy (mathmale):

right, and what is 9*9?

OpenStudy (anonymous):

why would u time 9 x 9??

OpenStudy (anonymous):

the fractions are 19/3 and 13/3?

OpenStudy (mathmale):

You're right; I'm wrong. Can you fix my mistake? What will the product now look like?

OpenStudy (anonymous):

19/3 x 13/3

OpenStudy (johnweldon1993):

Not quite... Change 1 5/8 into an improper fraction again... 13/8 is what I get there...so the fractions are \large \frac{19}{3} \times \frac{13}{8}\]

OpenStudy (mathmale):

\[\frac{ 19 }{ 3 }*\frac{ 13 }{ 3 }=?\]

OpenStudy (anonymous):

so fine 19 x 13 is 247 and then what?

OpenStudy (anonymous):

its 247/9 reduce down to 27.4

OpenStudy (anonymous):

yes thats the eqution

OpenStudy (mathmale):

Sorry, wrong again. Did anyone catch my mistake?

OpenStudy (johnweldon1993):

\[\large \frac{19}{3} \times \frac{13}{8}\]

OpenStudy (anonymous):

its 10 then

OpenStudy (mathmale):

Hint: Refer back to the original problem!! Kudos to @johonweldon1993!

OpenStudy (johnweldon1993):

You just improperly changed the 1 3/8 into an improper fraction...it should have been 13/8 not 13/3 :)

OpenStudy (mathmale):

Right! Thank you.

OpenStudy (whpalmer4):

3, 8, older eyes, small fonts, who can tell the difference? :-)

OpenStudy (anonymous):

life saver @johnweldon1993 lol

OpenStudy (anonymous):

ok but i have no clue how you got that when the problem was 6 1/3 times 1 5/8

OpenStudy (mathmale):

@c1282 : so, would you please do that multiplication?\[\frac{ 19 }{ 3 }*\frac{ 13 }{ 8 }=?\]

OpenStudy (johnweldon1993):

Alright...so to change a mixed number to an improper fraction... \[\large 1 \frac{5}{8}\] You take the denominator (8) and multiply it by the whole number (1)...then you add the numerator (5) to that result ... so that comes out to 13 ....but then you put it over the original denominator which was 8... Make sense @c1282 ?

OpenStudy (anonymous):

u multiply 6 times 3 plus 1 gives u 19 over three and 8 times 1 pplus 5 gives u 13 over 8

OpenStudy (anonymous):

i got that part... then i have 247 /24?

OpenStudy (whpalmer4):

Yes, that's correct.

OpenStudy (mathmale):

@gentleman313 : Nice work!

OpenStudy (johnweldon1993):

Correct

OpenStudy (mathmale):

Can 247/24 be reduced?

OpenStudy (anonymous):

divide 247 by 24

OpenStudy (anonymous):

then what divide the 247 and 24

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

Not quite divide..that will give you a decimal and as the question states...to leave in an improper fraction

OpenStudy (johnweldon1993):

Just reduce :)

OpenStudy (anonymous):

but it says it should be a whole number...

OpenStudy (mathmale):

Can you do that? 247/24, left as is, is the most accurate answer. 247/24, done on a calculator, results in only an approx. answer. Much depends upon the format in which your teacher wants your answers. Personally, I'd leave the answer as 247/24.

OpenStudy (whpalmer4):

"your answer should be a whole number, fraction or improper fraction only"

OpenStudy (anonymous):

well it needs to b ereduced and either a whole number improper or fraction

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh ok *facepalm

OpenStudy (johnweldon1993):

Right...whole number does not include decimals...247/24 would equate to 10.29 <-- not a whole number :)

OpenStudy (anonymous):

*Facepalm

OpenStudy (mathmale):

In that case, you'll need to round down to the nearest whole number. However... (a big However), your answer must be " either a whole number or improper fraction"...So, what 's YOUR final answer? Hint: be sure you know what an "improper fraction" is.

OpenStudy (anonymous):

i have no clue but the answer was 42 17/24????

OpenStudy (mathmale):

Would 6 x 2 come out to approx. 42? Estimate to judge whether your answer is reasonable or not.

OpenStudy (whpalmer4):

My comment was a direct quote from the problem statement: "our answer should be a whole number, fraction or improper fraction only" We don't have a whole number, but either a (proper) or improper fraction will suffice.

OpenStudy (mathmale):

So, @c1282 , what 's YOUR final answer? Hint: be sure you know what an "improper fraction" is.